OpenStudy (anonymous):

How many liters of chlorine gas, at standard temperature and pressure, will react with 15.5 grams of potassium metal? Show all of the work used to solve this problem. 2 K + Cl2 yields 2 KCl and How many liters of methane gas (CH4) need to be combusted to produce 12.4 liters of water vapor, if all measurements are taken at the same temperature and pressure? Show all of the work used to solve this problem. CH4 (g) + 2 O2 (g) yields CO2 (g) + 2 H2O (g)

OpenStudy (anonymous):

Oh I got the second problem done its: You know moles = volume/molar volume (Vm) n(H2O) = 12.4/(Vm) n(CH4)=n(H2O) * 0.5 = 6.2/Vm V(CH4) = 6.4/Vm *Vm = 6.2L

OpenStudy (aaronq):

"How many liters of chlorine gas, at standard temperature and pressure, will react with 15.5 grams of potassium metal? Show all of the work used to solve this problem. 2 K + Cl2 yields 2 KCl" Convert the mass of K into moles: \(\sf moles=\dfrac{mass}{Molar~mass}\) Use a ratio of the moles with the coefficients to find moles of \(Cl_2\) \(\sf \dfrac{moles~of~K}{K's~coefficient}=\dfrac{moles~of~Cl_2}{Cl_2's ~coefficient}\) Now find the L, 1 mole = 22.414 L